OpenStudy (anonymous):

What are the coordinates of the foci of the conic section shown below? (x+2)^2/16-(y-3)^2/9=1

OpenStudy (anonymous):

a. (-2±5,3) b. (-2,3±5) c. (2,3±√7) d. (-2±√7,3)

OpenStudy (anonymous):

ok here centre is at (-2,3) here a=4 b=3 c^2=a^2+b^2 \[\Large c=\sqrt{a^2+b^2}=\sqrt{16+9}=\sqrt{25}=5\] foci are (-c,0) and (+c,0) or \[\Large (\pm C,0)=(\pm 5,0)\] but center is at (-2,3) so \[\Large (-2 \pm 5,3)\] option a is more appropriate !

OpenStudy (anonymous):

did you get it ? @SantanaG

OpenStudy (anonymous):

yeah thank you! :)